# Sequential Conformal Risk Control UQ Uncertainty Quantification

### v1

@@ Appendix F Algorithms
mix box_count_threshold thresholded lac additive 17.7521 0.0179 10.9197 0.0484 12.5311 0.0476 0.0846
mix box_count_threshold thresholded lac multiplicative 17.7521 0.0179 6.4985 0.0512 12.5311 0.0476 0.0915
mix box_count_threshold thresholded aps additive 17.7521 0.0179 10.9197 0.0484 15.0931 0.0460 0.0857
mix box_count_threshold thresholded aps multiplicative 17.7521 0.0179 6.0228 0.0512 12.0499 0.0484 0.0923
mix box_count_threshold pixelwise lac additive 17.7521 0.0179 4.6034 0.0499 1.0064 0.0511 0.0871
mix box_count_threshold pixelwise lac multiplicative 17.7521 0.0179 2.8759 0.0512 1.0064 0.0511 0.0915
mix box_count_threshold pixelwise aps additive 17.7521 0.0179 4.6034 0.0499 1.0000 0.0491 0.0868
mix box_count_threshold pixelwise aps multiplicative 17.7521 0.0179 2.8759 0.0512 1.0000 0.0491 0.0915
mix box_count_threshold boxwise lac additive 17.7521 0.0179 5.1730 0.0499 1.0064 0.0511 0.0871
mix box_count_threshold boxwise lac multiplicative 17.7521 0.0179 3.6850 0.0512 1.0064 0.0511 0.0915
mix box_count_threshold boxwise aps additive 17.7521 0.0179 5.1730 0.0499 1.0000 0.0491 0.0868
mix box_count_threshold boxwise aps multiplicative 17.7521 0.0179 3.6850 0.0512 1.0000 0.0491 0.0915
mix box_count_recall thresholded lac additive 10.6552 0.0186 10.9197 0.0484 12.5311 0.0476 0.0846
mix box_count_recall thresholded lac multiplicative 10.6552 0.0186 6.4985 0.0512 12.5311 0.0476 0.0915
mix box_count_recall thresholded aps additive 10.6552 0.0186 10.9197 0.0484 15.0931 0.0460 0.0857
mix box_count_recall thresholded aps multiplicative 10.6552 0.0186 6.0228 0.0512 12.0499 0.0484 0.0923
mix box_count_recall pixelwise lac additive 10.6552 0.0186 4.6034 0.0499 1.0064 0.0511 0.0871
mix box_count_recall pixelwise lac multiplicative 10.6552 0.0186 2.8759 0.0512 1.0064 0.0511 0.0915
mix box_count_recall pixelwise aps additive 10.6552 0.0186 4.6034 0.0499 1.0000 0.0491 0.0868
mix box_count_recall pixelwise aps multiplicative 10.6552 0.0186 2.8759 0.0512 1.0000 0.0491 0.0915
mix box_count_recall boxwise lac additive 10.6552 0.0186 5.1730 0.0499 1.0064 0.0511 0.0871
mix box_count_recall boxwise lac multiplicative 10.6552 0.0186 3.6850 0.0512 1.0064 0.0511 0.0915
mix box_count_recall boxwise aps additive 10.6552 0.0186 5.1730 0.0499 1.0000 0.0491 0.0868
mix box_count_recall boxwise aps multiplicative 10.6552 0.0186 3.6850 0.0512 1.0000 0.0491 0.0915
giou box_count_threshold thresholded lac additive 17.7521 0.0179 26.2773 0.0476 35.0525 0.0484 0.0860
giou box_count_threshold thresholded lac multiplicative 17.7521 0.0179 20.1749 0.0499 35.0525 0.0484 0.0909
giou box_count_threshold thresholded aps additive 17.7521 0.0179 26.2773 0.0476 37.3819 0.0460 0.0853
giou box_count_threshold thresholded aps multiplicative 17.7521 0.0179 19.6941 0.0499 31.2916 0.0484 0.0909
giou box_count_threshold pixelwise lac additive 17.7521 0.0179 22.4286 0.0491 35.0525 0.0484 0.0868
giou box_count_threshold pixelwise lac multiplicative 17.7521 0.0179 19.1563 0.0509 35.0525 0.0484 0.0909
giou box_count_threshold pixelwise aps additive 17.7521 0.0179 22.4286 0.0491 37.3819 0.0460 0.0853
giou box_count_threshold pixelwise aps multiplicative 17.7521 0.0179 18.9911 0.0509 31.2916 0.0484 0.0909
giou box_count_threshold boxwise lac additive 17.7521 0.0179 23.3271 0.0491 35.0525 0.0484 0.0868
giou box_count_threshold boxwise lac multiplicative 17.7521 0.0179 20.1749 0.0499 35.0525 0.0484 0.0909
giou box_count_threshold boxwise aps additive 17.7521 0.0179 23.3271 0.0491 37.3819 0.0460 0.0853
giou box_count_threshold boxwise aps multiplicative 17.7521 0.0179 19.6941 0.0499 31.2916 0.0484 0.0909
giou box_count_recall thresholded lac additive 10.6552 0.0186 26.2773 0.0476 35.0525 0.0484 0.0860
giou box_count_recall thresholded lac multiplicative 10.6552 0.0186 20.1749 0.0499 35.0525 0.0484 0.0909
giou box_count_recall thresholded aps additive 10.6552 0.0186 26.2773 0.0476 37.3819 0.0460 0.0853
giou box_count_recall thresholded aps multiplicative 10.6552 0.0186 19.6941 0.0499 31.2916 0.0484 0.0909
giou box_count_recall pixelwise lac additive 10.6552 0.0186 22.4286 0.0491 35.0525 0.0484 0.0868
giou box_count_recall pixelwise lac multiplicative 10.6552 0.0186 19.1563 0.0509 35.0525 0.0484 0.0909
giou box_count_recall pixelwise aps additive 10.6552 0.0186 22.4286 0.0491 37.3819 0.0460 0.0853
giou box_count_recall pixelwise aps multiplicative 10.6552 0.0186 18.9911 0.0509 31.2916 0.0484 0.0909
giou box_count_recall boxwise lac additive 10.6552 0.0186 23.3271 0.0491 35.0525 0.0484 0.0868
giou box_count_recall boxwise lac multiplicative 10.6552 0.0186 20.1749 0.0499 35.0525 0.0484 0.0909
giou box_count_recall boxwise aps additive 10.6552 0.0186 23.3271 0.0491 37.3819 0.0460 0.0853
giou box_count_recall boxwise aps multiplicative 10.6552 0.0186 19.6941 0.0499 31.2916 0.0484 0.0909
hausdorff box_count_threshold thresholded lac additive 17.7521 0.0179 4.6034 0.0499 34.3152 0.0491 0.0883
hausdorff box_count_threshold thresholded lac multiplicative 17.7521 0.0179 2.9569 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_threshold thresholded aps additive 17.7521 0.0179 4.6034 0.0499 32.8959 0.0476 0.0879
hausdorff box_count_threshold thresholded aps multiplicative 17.7521 0.0179 2.9569 0.0512 32.8959 0.0476 0.0915
hausdorff box_count_threshold pixelwise lac additive 17.7521 0.0179 4.6034 0.0499 34.3152 0.0491 0.0883
hausdorff box_count_threshold pixelwise lac multiplicative 17.7521 0.0179 2.5616 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_threshold pixelwise aps additive 17.7521 0.0179 4.6034 0.0499 32.8959 0.0476 0.0879
hausdorff box_count_threshold pixelwise aps multiplicative 17.7521 0.0179 2.5616 0.0512 32.8959 0.0476 0.0915
hausdorff box_count_threshold boxwise lac additive 17.7521 0.0179 3.9352 0.0510 34.3152 0.0491 0.0887
hausdorff box_count_threshold boxwise lac multiplicative 17.7521 0.0179 2.9569 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_threshold boxwise aps additive 17.7521 0.0179 3.9352 0.0510 32.8959 0.0476 0.0879
hausdorff box_count_threshold boxwise aps multiplicative 17.7521 0.0179 2.9569 0.0512 32.8959 0.0476 0.0915
hausdorff box_count_recall thresholded lac additive 10.6552 0.0186 4.6034 0.0499 34.3152 0.0491 0.0883
hausdorff box_count_recall thresholded lac multiplicative 10.6552 0.0186 2.9569 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_recall thresholded aps additive 10.6552 0.0186 4.6034 0.0499 32.8959 0.0476 0.0879
hausdorff box_count_recall thresholded aps multiplicative 10.6552 0.0186 2.9569 0.0512 32.8959 0.0476 0.0915
hausdorff box_count_recall pixelwise lac additive 10.6552 0.0186 4.6034 0.0499 34.3152 0.0491 0.0883
hausdorff box_count_recall pixelwise lac multiplicative 10.6552 0.0186 2.5616 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_recall pixelwise aps additive 10.6552 0.0186 4.6034 0.0499 32.8959 0.0476 0.0879
hausdorff box_count_recall pixelwise aps multiplicative 10.6552 0.0186 2.5616 0.0512 32.8959 0.0476 0.0915
hausdorff box_count_recall boxwise lac additive 10.6552 0.0186 3.9352 0.0510 34.3152 0.0491 0.0887
hausdorff box_count_recall boxwise lac multiplicative 10.6552 0.0186 2.9569 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_recall boxwise aps additive 10.6552 0.0186 3.9352 0.0510 32.8959 0.0476 0.0879
hausdorff box_count_recall boxwise aps multiplicative 10.6552 0.0186 2.9569 0.0512 32.8959 0.0476 0.0915
lac box_count_threshold thresholded lac additive 17.7521 0.0179 18.2558 0.0491 1.0064 0.0511 0.0868
lac box_count_threshold thresholded lac multiplicative 17.7521 0.0179 12.1580 0.0509 1.0064 0.0511 0.0909
lac box_count_threshold thresholded aps additive 17.7521 0.0179 18.2558 0.0491 1.0000 0.0491 0.0864
lac box_count_threshold thresholded aps multiplicative 17.7521 0.0179 12.1580 0.0509 1.0000 0.0491 0.0909
lac box_count_threshold pixelwise lac additive 17.7521 0.0179 14.4720 0.0499 1.0064 0.0511 0.0871
lac box_count_threshold pixelwise lac multiplicative 17.7521 0.0179 10.8107 0.0512 1.0064 0.0511 0.0915
lac box_count_threshold pixelwise aps additive 17.7521 0.0179 14.4720 0.0499 1.0000 0.0491 0.0868
lac box_count_threshold pixelwise aps multiplicative 17.7521 0.0179 10.8107 0.0512 1.0000 0.0491 0.0915
lac box_count_threshold boxwise lac additive 17.7521 0.0179 15.8677 0.0499 1.0064 0.0511 0.0871
lac box_count_threshold boxwise lac multiplicative 17.7521 0.0179 12.1580 0.0509 1.0064 0.0511 0.0909
lac box_count_threshold boxwise aps additive 17.7521 0.0179 15.8677 0.0499 1.0000 0.0491 0.0868
lac box_count_threshold boxwise aps multiplicative 17.7521 0.0179 12.1580 0.0509 1.0000 0.0491 0.0909
lac box_count_recall thresholded lac additive 10.6552 0.0186 18.2558 0.0491 1.0064 0.0511 0.0868
lac box_count_recall thresholded lac multiplicative 10.6552 0.0186 12.1580 0.0509 1.0064 0.0511 0.0909
lac box_count_recall thresholded aps additive 10.6552 0.0186 18.2558 0.0491 1.0000 0.0491 0.0864
lac box_count_recall thresholded aps multiplicative 10.6552 0.0186 12.1580 0.0509 1.0000 0.0491 0.0909
lac box_count_recall pixelwise lac additive 10.6552 0.0186 14.4720 0.0499 1.0064 0.0511 0.0871
lac box_count_recall pixelwise lac multiplicative 10.6552 0.0186 10.8107 0.0512 1.0064 0.0511 0.0915
lac box_count_recall pixelwise aps additive 10.6552 0.0186 14.4720 0.0499 1.0000 0.0491 0.0868
lac box_count_recall pixelwise aps multiplicative 10.6552 0.0186 10.8107 0.0512 1.0000 0.0491 0.0915
lac box_count_recall boxwise lac additive 10.6552 0.0186 15.8677 0.0499 1.0064 0.0511 0.0871
lac box_count_recall boxwise lac multiplicative 10.6552 0.0186 12.1580 0.0509 1.0064 0.0511 0.0909
lac box_count_recall boxwise aps additive 10.6552 0.0186 15.8677 0.0499 1.0000 0.0491 0.0868
lac box_count_recall boxwise aps multiplicative 10.6552 0.0186 12.1580 0.0509 1.0000 0.0491 0.0909

### v2

@@ Appendix F Effect of Object Detection and Conformal Settings
mix box_count_threshold thresholded lac additive 18.8552 0.0124 6.2473 0.0464 69.8821 0.0517 0.0943
mix box_count_threshold thresholded lac multiplicative 18.8552 0.0124 6.6374 0.052 69.8821 0.0517 0.1021
mix box_count_threshold thresholded aps additive 18.8552 0.0124 6.2473 0.0464 40.7516 0.0418 0.0852
mix box_count_threshold thresholded aps multiplicative 18.8552 0.0124 6.6374 0.052 40.7516 0.0418 0.0929
mix box_count_threshold pixelwise lac additive 18.8552 0.0124 4.3435 0.0477 69.8823 0.0517 0.092
mix box_count_threshold pixelwise lac multiplicative 18.8552 0.0124 4.1814 0.0517 69.8823 0.0517 0.098
mix box_count_threshold pixelwise aps additive 18.8552 0.0124 4.3435 0.0477 40.7516 0.0418 0.0816
mix box_count_threshold pixelwise aps multiplicative 18.8552 0.0124 4.1814 0.0517 40.7516 0.0418 0.0885
mix box_count_threshold boxwise lac additive 18.8552 0.0124 6.1211 0.0432 69.8823 0.0517 0.0883
mix box_count_threshold boxwise lac multiplicative 18.8552 0.0124 5.8593 0.0495 69.8823 0.0517 0.0961
mix box_count_threshold boxwise aps additive 18.8552 0.0124 6.1211 0.0432 40.7516 0.0418 0.0781
mix box_count_threshold boxwise aps multiplicative 18.8552 0.0124 5.8593 0.0495 40.7516 0.0418 0.0869
mix box_count_recall thresholded lac additive 11.7104 0.0189 5.4582 0.0456 67.5809 0.0531 0.0936
mix box_count_recall thresholded lac multiplicative 11.7104 0.0189 6.961 0.0544 67.5809 0.0531 0.1061
mix box_count_recall thresholded aps additive 11.7104 0.0189 5.4582 0.0456 41.6644 0.0416 0.0835
mix box_count_recall thresholded aps multiplicative 11.7104 0.0189 6.961 0.0544 41.6644 0.0416 0.0948
mix box_count_recall pixelwise lac additive 11.7104 0.0189 3.867 0.0487 67.5809 0.0531 0.0935
mix box_count_recall pixelwise lac multiplicative 11.7104 0.0189 4.5352 0.0492 67.5809 0.0531 0.0976
mix box_count_recall pixelwise aps additive 11.7104 0.0189 3.867 0.0487 41.6644 0.0416 0.0814
mix box_count_recall pixelwise aps multiplicative 11.7104 0.0189 4.5352 0.0492 41.6644 0.0416 0.0859
mix box_count_recall boxwise lac additive 11.7104 0.0189 5.2321 0.0453 67.5809 0.0531 0.0902
mix box_count_recall boxwise lac multiplicative 11.7104 0.0189 6.2333 0.0477 67.5809 0.0531 0.0965
mix box_count_recall boxwise aps additive 11.7104 0.0189 5.2321 0.0453 41.6644 0.0416 0.0791
mix box_count_recall boxwise aps multiplicative 11.7104 0.0189 6.2333 0.0477 41.6644 0.0416 0.085
hausdorff box_count_threshold thresholded lac additive 18.8552 0.0124 6.2473 0.046 70.2243 0.0523 0.0945
hausdorff box_count_threshold thresholded lac multiplicative 18.8552 0.0124 7.4398 0.0464 70.2243 0.0523 0.0975
hausdorff box_count_threshold thresholded aps additive 18.8552 0.0124 6.2473 0.046 40.9013 0.0434 0.0864
hausdorff box_count_threshold thresholded aps multiplicative 18.8552 0.0124 7.4398 0.0464 40.9012 0.0434 0.089
hausdorff box_count_threshold pixelwise lac additive 18.8552 0.0124 4.3435 0.0477 70.2243 0.0523 0.0926
hausdorff box_count_threshold pixelwise lac multiplicative 18.8552 0.0124 4.6191 0.0473 70.2243 0.0523 0.0952
hausdorff box_count_threshold pixelwise aps additive 18.8552 0.0124 4.3435 0.0477 40.9013 0.0434 0.0829
hausdorff box_count_threshold pixelwise aps multiplicative 18.8552 0.0124 4.6191 0.0473 40.9013 0.0434 0.0862
hausdorff box_count_threshold boxwise lac additive 18.8552 0.0124 6.1211 0.0431 70.2243 0.0523 0.0886
hausdorff box_count_threshold boxwise lac multiplicative 18.8552 0.0124 6.5888 0.0452 70.2243 0.0523 0.0934
hausdorff box_count_threshold boxwise aps additive 18.8552 0.0124 6.1211 0.0431 40.9013 0.0434 0.0796
hausdorff box_count_threshold boxwise aps multiplicative 18.8552 0.0124 6.5888 0.0452 40.9012 0.0434 0.0846
hausdorff box_count_recall thresholded lac additive 11.7104 0.0189 5.4513 0.0468 67.9788 0.0539 0.0957
hausdorff box_count_recall thresholded lac multiplicative 11.7104 0.0189 8.0526 0.044 67.9788 0.0539 0.0967
hausdorff box_count_recall thresholded aps additive 11.7104 0.0189 5.4513 0.0468 41.7312 0.0428 0.0858
hausdorff box_count_recall thresholded aps multiplicative 11.7104 0.0189 8.0526 0.044 41.7312 0.0428 0.0857
hausdorff box_count_recall pixelwise lac additive 11.7104 0.0189 3.867 0.0487 67.9788 0.0539 0.0945
hausdorff box_count_recall pixelwise lac multiplicative 11.7104 0.0189 4.9961 0.0438 67.9788 0.0539 0.0939
hausdorff box_count_recall pixelwise aps additive 11.7104 0.0189 3.867 0.0487 41.7312 0.0428 0.0824
hausdorff box_count_recall pixelwise aps multiplicative 11.7104 0.0189 4.9961 0.0438 41.7312 0.0428 0.0821
hausdorff box_count_recall boxwise lac additive 11.7104 0.0189 5.2252 0.0457 67.9788 0.0539 0.0918
hausdorff box_count_recall boxwise lac multiplicative 11.7104 0.0189 6.9853 0.0416 67.9788 0.0539 0.0921
hausdorff box_count_recall boxwise aps additive 11.7104 0.0189 5.2252 0.0457 41.7312 0.0428 0.0808
hausdorff box_count_recall boxwise aps multiplicative 11.7104 0.0189 6.9853 0.0416 41.7312 0.0428 0.0807
lac box_count_threshold thresholded lac additive 18.8552 0.0124 25.0369 0.0504 0.7168 0.0489 0.0946
lac box_count_threshold thresholded lac multiplicative 18.8552 0.0124 21.6649 0.0428 0.7168 0.0489 0.0869
lac box_count_threshold thresholded aps additive 18.8552 0.0124 25.0369 0.0504 0.996 0.0428 0.0893
lac box_count_threshold thresholded aps multiplicative 18.8552 0.0124 21.6649 0.0428 0.996 0.0428 0.0817
lac box_count_threshold pixelwise lac additive 18.8552 0.0124 20.6405 0.0467 0.7168 0.0489 0.086
lac box_count_threshold pixelwise lac multiplicative 18.8552 0.0124 15.8047 0.0491 0.7168 0.0489 0.0853
lac box_count_threshold pixelwise aps additive 18.8552 0.0124 20.6405 0.0467 0.996 0.0428 0.0817
lac box_count_threshold pixelwise aps multiplicative 18.8552 0.0124 15.8047 0.0491 0.996 0.0428 0.0812
lac box_count_threshold boxwise lac additive 18.8552 0.0124 24.1811 0.0459 0.7168 0.0489 0.0828
lac box_count_threshold boxwise lac multiplicative 18.8552 0.0124 17.8473 0.0486 0.7168 0.0489 0.0838
lac box_count_threshold boxwise aps additive 18.8552 0.0124 24.1811 0.0459 0.996 0.0428 0.079
lac box_count_threshold boxwise aps multiplicative 18.8552 0.0124 17.8473 0.0486 0.996 0.0428 0.0797
lac box_count_recall thresholded lac additive 11.7104 0.0189 21.8857 0.0388 1.0285 0.0489 0.083
lac box_count_recall thresholded lac multiplicative 11.7104 0.0189 25.4212 0.0348 1.0285 0.0489 0.079
lac box_count_recall thresholded aps additive 11.7104 0.0189 21.8857 0.0388 1.3236 0.0478 0.0822
lac box_count_recall thresholded aps multiplicative 11.7104 0.0189 25.4212 0.0348 1.3236 0.0478 0.0786
lac box_count_recall pixelwise lac additive 11.7104 0.0189 17.6658 0.0437 1.0285 0.0489 0.0823
lac box_count_recall pixelwise lac multiplicative 11.7104 0.0189 17.7072 0.0426 1.0285 0.0489 0.0792
lac box_count_recall pixelwise aps additive 11.7104 0.0189 17.6658 0.0437 1.3236 0.0478 0.0822
lac box_count_recall pixelwise aps multiplicative 11.7104 0.0189 17.7072 0.0426 1.3236 0.0478 0.0795
lac box_count_recall boxwise lac additive 11.7104 0.0189 20.4723 0.0425 1.0285 0.0489 0.0797
lac box_count_recall boxwise lac multiplicative 11.7104 0.0189 20.036 0.0411 1.0285 0.0489 0.077
lac box_count_recall boxwise aps additive 11.7104 0.0189 20.4723 0.0425 1.3236 0.0478 0.0796
lac box_count_recall boxwise aps multiplicative 11.7104 0.0189 20.036 0.0411 1.3236 0.0478 0.0772
giou box_count_threshold thresholded lac additive 18.8552 0.0124 43.6338 0.016 65.9061 0.0401 0.0559
giou box_count_threshold thresholded aps additive 18.8552 0.0124 43.6338 0.016 65.4388 0.0268 0.0421
giou box_count_threshold pixelwise lac additive 18.8552 0.0124 39.4903 0.0258 65.906 0.0401 0.0655
giou box_count_threshold pixelwise lac multiplicative 18.8552 0.0124 171.0406 0.0154 65.906 0.0401 0.0553
giou box_count_threshold pixelwise aps additive 18.8552 0.0124 39.4903 0.0258 65.4388 0.0268 0.0506
giou box_count_threshold pixelwise aps multiplicative 18.8552 0.0124 171.0406 0.0154 65.4388 0.0268 0.0417
giou box_count_threshold boxwise lac additive 18.8552 0.0124 43.3738 0.0198 65.906 0.0401 0.0595
giou box_count_threshold boxwise lac multiplicative 18.8552 0.0124 195.5759 0.0136 65.906 0.0401 0.0534
giou box_count_threshold boxwise aps additive 18.8552 0.0124 43.3738 0.0198 65.4388 0.0268 0.0448
giou box_count_threshold boxwise aps multiplicative 18.8552 0.0124 195.5759 0.0136 65.4388 0.0268 0.0399
giou box_count_recall thresholded lac additive 11.7104 0.0189 35.7372 0.0092 67.048 0.0417 0.0504
giou box_count_recall thresholded aps additive 11.7104 0.0189 35.7372 0.0092 66.5712 0.024 0.0325
giou box_count_recall pixelwise lac additive 11.7104 0.0189 32.3787 0.0192 67.0478 0.0417 0.0602
giou box_count_recall pixelwise lac multiplicative 11.7104 0.0189 171.6958 0.0073 67.0478 0.0417 0.0486
giou box_count_recall pixelwise aps additive 11.7104 0.0189 32.3787 0.0192 66.5712 0.024 0.041
giou box_count_recall pixelwise aps multiplicative 11.7104 0.0189 171.6958 0.0073 66.5712 0.024 0.0306
giou box_count_recall boxwise lac additive 11.7104 0.0189 35.5256 0.0128 67.0478 0.0417 0.0538
giou box_count_recall boxwise lac multiplicative 11.7104 0.0189 196.7056 0.0067 67.0478 0.0417 0.048
giou box_count_recall boxwise aps additive 11.7104 0.0189 35.5256 0.0128 66.5712 0.024 0.0349
giou box_count_recall boxwise aps multiplicative 11.7104 0.0189 196.7056 0.0067 66.5712 0.024 0.0299